# ATOMIUM: Dust and tracers of binarity in the continua

### v1

@@ 5.4 Two modes of dust formation: pulsation- and binary-enhanced
We have identified two pathways for shock-driven dust formation: stellar pulsations and the supersonic motion of companion stars through the circumstellar envelope. All stars in our sample exhibit some dust in the close stellar environment (even U Del, see Sect. 5.1.5). However, not all stars exhibit extended regions of continuum flux. Excluding RW Sco101010RW Sco is moderately distant (560 pc) and, in general, has a lower signal to noise than the majority of our sample, so it is possible that our observations were not sensitive enough to detect or resolve regions of extended flux., all the stars with periods longer than 300 days have some extended regions of continuum flux. For periods shorter than 300 days, only two unusual stars have notable regions of extended flux: Gru and SV Aqr, discussed above in Sections 5.1.1 and 5.1.4, respectively. In both of these cases the extended dust emission can be attributed to a known ( Gru) or suspected (SV Aqr, see Appendix A) companion.
@@ 5.4.1 Pulsation-enhanced dust formation
Pulsation periods of days are thought to be necessary for an AGB star to fully develop a pulsation-enhanced dust-driven wind (Winters et al., 2000; McDonald & Zijlstra, 2016), a hypothesis supported by observations of both oxygen-rich and carbon-rich stars, where a positive correlation between mass-loss rate and period is seen above 300 days (Groenewegen et al., 1998; Olofsson et al., 2002; Ramstedt et al., 2009; McDonald & Zijlstra, 2016). This is also in agreement with the relation found by Wood (1990) for an exponential increase in mass-loss rate with pulsation period111111The relation found by Wood (1990) is , mainly holding between 300 and 500 day periods, above which the mass-loss rate no longer increases with period (Vassiliadis & Wood, 1993).
@@
McDonald & Zijlstra (2016) also noted that stars with pulsation periods above 500 days were the most extreme stars in the sample and included RSGs and known binaries. When considering the period–radius and period–monochromatic luminosity plots in Fig. 5, we see an uptick in radius among the Miras for periods above 300 days, possibly suggesting that longer pulsation periods correlate to larger radii as well as increased dust production for this period range. However, the uptick in monochromatic luminosity is only seen above day periods, suggesting that it is only these most extreme stars which have higher mm luminosities. This could possibly be the result of the UD fits to the 5 extreme stars (two extreme and distant AGB stars and our RSG subsample) including more circumstellar dust than the closer stars in our sample. This is partly reflected in the rather low brightness temperatures obtained for IRC10529 and AH Sco. Kiss et al. (2006) also noted that VX Sgr (and to a lesser extent AH Sco121212Note also that the Kiss et al. (2006) sample did not include KW Sgr.) has a particularly large pulsation amplitude, which they associated with increased pulsation-enhanced mass loss and dust production. This fits with the finding by Freytag & Höfner (2023), based on their 3D theoretical models, that dust forms in the wake of shocks within a few stellar radii of the star; larger-amplitude pulsations would produce stronger shocks. All three RSGs in ATOMIUM agree well with a trend with pulsation period, even though KW Sgr appears to be the least dusty of the three. IRC10529 and IRC+10011 are also known to be extremely dusty (Justtanont et al., 2013; Reiter et al., 2015), which fits with the “extreme” label noted by McDonald & Zijlstra (2016). From these observations we can hypothesise that for stars with periods above 300 days we see an increase in both mass-loss rate and mm radius, while for stars with periods above 500 days we see an increase in monochromatic luminosity and another increase in mass-loss rate.
@@ 5.4.2 Binary-enhanced dust formation
Gru and R Aql have the two highest mass-loss rates for stars with periods below 300 days. For R Aql this can be explained by its declining period (Zhao-Geisler et al., 2012); in 1915 the period was days, placing it within the trend of the other longer-period stars. The mass-loss rate calculated from CO lines is an average over several hundred years so will not have altered appreciably in that time. For Gru, the mass-loss rate calculated by Doan et al. (2017) takes the complex circumstellar structure into account. The high mass-loss rate, compared with other stars with pulsation periods around 200 days, can be explained by the extra dust produced in the wake of the companion, which orbits with a semi-major axis of au (approx. 3–4 stellar radii, Montargès et al. 2025, Esseldeurs et al, subm.). Once more dust is formed, radiation pressure from the AGB star will push it outwards with the rest of the dust. If it is within the wind acceleration region, this additional dust can enhance the mass loss as part of the dust-driven wind, since there will be more dust available to collide with and drag the gas outwards.
@@
The 3D single star models of Freytag & Höfner (2023) show that some gas still falls back onto the star out to radii of au, so additional dust produced within 9 au could contribute to enhanced mass-loss. This phenomenon could explain moderate mass-loss rates for the majority of surveyed SR stars with shorter periods (¡300 days). Because the shocks induced by the companion persist further out in the wind (e.g. in spiral patterns, see Appendix F and e.g. Kim et al. 2017, 2019, Malfait et al. 2021), some dust will also form well outside of the wind acceleration region (see the middle panel of Fig. 20), and will increase the overall dust-to-gas ratio compared with that of an otherwise similar single star. This could also contribute to the over-estimation of mass-loss rates calculated based on dust properties and assumed dust-to-gas ratios and could explain the discrepancy in mass-loss rates calculated from dust and from CO for the dustiest stars (e.g. see comparisons in Table 4 of Justtanont et al., 2006, where mass-loss rates from the two methods differ by around an order of magnitude).
@@
Many surveys, including both ATOMIUM and SUCCESS, are biased towards stars that have previously been detected (strongly) in CO, meaning they are likely to have higher mass-loss rates. If single stars with short pulsation periods have low levels of dust production and low mass-loss rates (e.g. McDonald et al., 2018) they would be less likely to be included in biased surveys. This explains the apparently high prevalence of binary AGB stars judging by their CO morphologies in high resolution ALMA observations. An unbiased survey such as the Nearby Evolved Star Survey (NESS, Scicluna et al., 2022) is hence likely to include stars with much lower mass-loss rates than stars that have been more frequently studied for their brighter emission lines. Another example is the small fraction of dusty field stars with luminosities below the RGB tip (McDonald et al., 2012, 2017). These have been assumed to be the lowest-mass AGB stars, since they have similar luminosities to the lowest-luminosity dusty AGB stars in globular clusters (Boyer et al., 2009; McDonald et al., 2009, 2011). However, these could instead be stars that have dust production triggered by companions. The corollary to this would be that the AGB stars with the lowest mass-loss rates (perhaps or lower) are least likely to have close companions.
@@
Another interesting example is the dusty but relatively low mass-loss rate AGB star L2 Pup (the lowest mass-loss rate source in Fig. 8). Kervella et al. (2016) found a low mass of for the companion, which is embedded in a dusty, nearly edge-on disc (Kervella et al., 2015; Ohnaka et al., 2015). The dust around L2 Pup is not distributed spherically, suggesting that the companion has strongly influenced its distribution. However, with a low mass, the gravitational potential of the companion is low, and it will cause weaker shocks as it orbits the AGB star (Maes et al., 2021), contributing to less dust formation than if it were more massive. This explains why the mass-loss rate of L2 Pup remains low (, Danilovich et al. 2015b and Appendix G).
At , around half of stars are expected to have companions with masses at least 10% that of the primary (Moe & Di Stefano, 2017). The proportion of stars with such companions increases to over 100% for stars of , meaning that some such stars are expected to be in triple systems. Close companions are more likely for higher-mass stars (Moe & Di Stefano, 2017). However, the closer the companion, the more intense the shockwaves generated as it travels through the wind and hence the more dust formation triggered, in addition to the pulsation-driven dust formation. W Aql B suffers from mag of extinction au from the AGB star (Danilovich et al., 2015a) but can still be detected with Hubble and SPHERE (Ramstedt et al., 2011; Montargès et al., 2023). The close companion to Gru has only been detected by the dust associated with it in the ATOMIUM continuum (Fig. LABEL:fig:pi1gru and Homan et al. 2020) and in scattered polarised light (Montargès et al., 2023, 2025). The optical light (whether the companion is a white dwarf or a main sequence star) has been completely attenuated by the dust.
Gru is the second closest star in the ATOMIUM sample and the companion is only just resolved from the AGB continuum in the extended data. If the system were a little further away, or oriented less favourably, or the projected separation between the two stars a little smaller, then we would not be able to distinguish the companion from the AGB continuum flux. Therefore, we cannot rule out the presence of companions for the majority of the ATOMIUM sample just because we have not (yet) directly observed them. Unseen companions could explain the extended dust distributions seen for U Her, R Hya, and S Pav, as well as the less symmetric emission seen for GY Aql, IRC10529, SV Aqr, T Mic and IRC+10011. This leaves R Aql, U Del, V PsA and RW Sco as the only AGB stars in our sample without evidence of binarity from their continuum properties.
@@ 5.5 Other indications of binarity
Indirect evidence of AGB binarity has previously been found in resolved CO emission, including from the ATOMIUM data (Decin et al., 2020) and earlier studies such as Maercker et al. (2012), Cernicharo et al. (2015) and Kim et al. (2017). Such structures can be reproduced — or partially reproduced — through hydrodynamic modelling (e.g. Mastrodemos & Morris, 1998, 1999; Kim et al., 2019; El Mellah et al., 2020; Malfait et al., 2021). Molecular signatures can also be used to indirectly infer the presence of UV-emitting binaries (Van de Sande & Millar, 2022; Siebert et al., 2022; Danilovich et al., 2024). Studies surveying AGB stars in the UV or X-ray have detected emission for some stars (e.g. Ramstedt et al., 2012; Sahai et al., 2015; Ortiz & Guerrero, 2021; Schmitt et al., 2024) which ought not to come from the AGB star itself. This emission is thought to arise from companion stars, most likely from accretion discs that have formed around the companions (Sahai et al., 2015). None of the stars in the ATOMIUM sample are included in the aforementioned studies as confirmed detections. Schmitt et al. (2024) include three ATOMIUM stars ( Gru, R Hya and S Pav) in their list of surveyed stars, but note optical contamination, which cannot be disentangled from potential X-ray detections.
@@
Out of the 14 AGB stars in the ATOMIUM sample, we have suggested that 10 of them might have binary companions, 8 of which have not been directly detected. Given that the ATOMIUM sample was selected to cover a range of mass-loss rates and pulsation behaviours (Gottlieb et al., 2022), there was no particular expectation for the stars to have binary companions. W Aql, Gru and R Hya were already know to have wide binary companions (see Sect. 2.1) and Gru has been found to be a triple system, which was suspected prior to the ATOMIUM observations (e.g. Chiu et al. 2006). Population statistics indicate that at least half of stars in the AGB mass range should have companions (Moe & Di Stefano, 2017), so earlier assumptions of the majority of AGB stars being single stars (i.e. before the advent of resolved imaging of the circumstellar environments) are statistically unlikely.
@@
Recent data and developments from Gaia have allowed us to better constrain distances and infer the presence of companions (including some which can be resolved Kervella et al., 2022). Based on the results of Kervella et al. (2022), Montargès et al. (2023) used a proper motion analysis combining Gaia Early Data Release 3 and Hipparcos data to estimate the probability of detecting close companions around the AGB stars. Note that the proper motion anomaly is not sensitive to wide binaries. They determine that R Hya, Gru, R Aql, and GY Aql are likely to have a detectable close companion based on the large S/N of the proper motion anomaly. Aside from R Aql, these agree with our predictions.
Another consideration is the link between dust and AGB mass loss. The concept of the dust-driven wind around AGB stars has long been the accepted mechanism for mass loss (Höfner & Olofsson, 2018). As discussed in Sect. 5.4, a close companion could amplify the amount of dust formed around the AGB star. This dust could then contribute to driving the wind, if it were formed sufficiently close to the AGB star. This would imply that AGB stars with close companions (i.e. within 10 stellar radii), could have higher mass-loss rates than those which do not. This does appear to be the case for Gru, which sits with a mass-loss rate above most of the other SR variables in Fig. 8. For W Aql, the dust forming in the wake of the companion au from the AGB star is unlikely to contribute to driving the wind. However, when the companion passed close to the AGB star during their periastron interaction (within 13 au, see Danilovich et al., 2024), it would have temporarily had a stronger effect on the dust formation.
Companion-enhanced dust formation could also explain the high mass-loss rates of the more extreme Mira variables in our sample: GY Aql, IRC10529 and IRC+10011. We also note that the carbon star with the highest mass-loss rate in the SUCCESS sample, comparable to the aforementioned oxygen-rich stars, is II Lup, around which Unnikrishnan et al. (2024) recently tentatively identified SiN, which suggests the presence of a binary companion (Van de Sande & Millar, 2022). The oxygen-rich star with the highest mass-loss rate in the SUCCESS sample is GX Mon, around which earlier observations have shown a circumstellar spiral, also indicative of a binary companion (Randall et al., 2020).
@@ 6 Summary and conclusions
We present the continuum data observed for a sample of 14 AGB and 3 RSG stars as part of the ATOMIUM ALMA Large Programme. For each star, we show the continuum observations from each individual array configuration as well as the combined image. Different features are apparent on different scales, owing to resolution, maximum recoverable scales and noise levels. We identify stars with extended continuum emission, bipolar structures and a few unique structures. Notably, we see a tail of dust in the wake of the close companion to Gru and some dust forming in the wake of the distant companion to W Aql. Out of 14 AGB stars in our sample, we see some evidence of binarity in the continuum images of 10 of them.
In addition to the main ALMA array data, we present new observations taken at lower frequencies and lower resolution with the ACA, for a subsample of 7 AGB stars. For these stars we calculate spectral indices, which range from 1.5–2.5, in agreement with theoretical expectations.
For each star, we fit uniform discs to the stellar emission and compare the resulting sizes and flux densities with other stellar properties such as pulsation period. We also present residual images after subtracting the uniform discs from the combined continuum maps. Among the AGB stars, we find that all the semiregular variables in our sample have lower monochromatic luminosities and smaller stellar radii than the Mira variables. As expected, the RSGs have even larger radii and higher monochromatic luminosities than the Mira variables.
@@
Collecting mass-loss rates from the literature for the ATOMIUM sample, we find that there is no correlation between pulsation period and mass-loss rate for periods below 300 days, and a positive trend above 300 days, as has been previously seen. Comparing the ATOMIUM sample to the larger SUCCESS survey, we find that among semiregular stars with low periods, Gru has an especially large mass-loss rate, and conclude this is likely because of the additional dust forming in the wake of the shocks generated by the passage of its close companion. Hence, we propose two channels for dust formation around AGB stars: in the wake of shocks caused by pulsations and in the wake of shocks caused by the supersonic passage of a companion through the circumstellar envelope. Consequently, we propose that extended dust emission or structure at mm wavelengths is an indirect indicator of potential binarity.

### v2

@@ 5.3 Two modes of dust formation: pulsation- and binary-enhanced
We propose a second pathway for shock-driven dust formation. In addition to the well established pathway of stellar pulsations, we propose that the supersonic motion of companion stars through the circumstellar envelope enhances dust production. All stars in our sample exhibit some dust in the close stellar environment (even U Del, on the basis of its long secondary period see Sect. 5.1.5). However, not all stars exhibit regions of continuum flux offset from the continuum peak. Excluding RW Sco121212RW Sco is moderately distant (560 pc) and, in general, has a lower signal to noise than the majority of our sample, so it is possible that our observations were not sensitive enough to detect or resolve regions of extended flux., all the stars with periods longer than 300 days have some resolved regions of continuum flux. For periods shorter than 300 days, only two unusual stars have notable regions of resolved flux: Gru and SV Aqr, discussed above in Sections 5.1.1 and 5.1.4, respectively. In both of these cases the dust emission offset from the continuum peak can be attributed to a known ( Gru) or suspected (SV Aqr, see Appendix A.1) companion.
@@ 5.3.1 Pulsation-enhanced dust formation
Pulsation periods of days are thought to be necessary for an AGB star to fully develop a pulsation-enhanced dust-driven wind (Winters et al., 2000; McDonald & Zijlstra, 2016), a hypothesis supported by observations of both oxygen-rich and carbon-rich stars, where a positive correlation between mass-loss rate and period is seen above 300 days (Groenewegen et al., 1998; Olofsson et al., 2002; Ramstedt et al., 2009; McDonald & Zijlstra, 2016). This is also in agreement with the relation found by Wood (1990) for an exponential increase in mass-loss rate with pulsation period131313The relation found by Wood (1990) is , mainly holding between 300 and 500 day periods, above which the mass-loss rate no longer increases with period (Vassiliadis & Wood, 1993).
@@
McDonald & Zijlstra (2016) also noted that stars with pulsation periods above 500 days were the most extreme stars in the sample and included RSGs and known binaries. When considering the period–radius and period–monochromatic luminosity plots in Fig. 7, we see an uptick in radius among the Miras for periods above 300 days, possibly suggesting that longer pulsation periods correlate to larger radii as well as increased dust production for this period range. However, the uptick in monochromatic luminosity is only seen above day periods, suggesting that it is only these most extreme stars which have higher mm luminosities. This could possibly be the result of the UD fits to the 5 extreme stars (two extreme and distant AGB stars and our RSG subsample) including more circumstellar dust than the closer stars in our sample. This is partly reflected in the rather low brightness temperatures obtained for IRC10529 and AH Sco. Kiss et al. (2006) also noted that VX Sgr (and to a lesser extent AH Sco141414Note also that the Kiss et al. (2006) sample did not include KW Sgr.) has a particularly large pulsation amplitude, which they associated with increased pulsation-enhanced mass loss and dust production. This fits with the finding by Freytag & Höfner (2023), based on their 3D theoretical models, that dust forms in the wake of shocks within a few stellar radii of the star; larger-amplitude pulsations would produce stronger shocks. All three RSGs in ATOMIUM agree well with a trend with pulsation period, even though KW Sgr appears to be the least dusty of the three (see estimated dust mass in Table 3). IRC10529 and IRC+10011 are also known to be extremely dusty (Justtanont et al., 2013; Reiter et al., 2015), which fits with the “extreme” label noted by McDonald & Zijlstra (2016). From these observations we can hypothesise that for stars with periods above 300 days we see an increase in both mass-loss rate and mm radius, while for stars with periods above 500 days we see an increase in monochromatic luminosity and another increase in mass-loss rate.
To check this hypothesis, we also plotted mass-loss rates for the AGB stars against each of UD radius and monochromatic luminosity, as shown in Fig. 11. In both plots, the two more extreme sources, IRC10529 and IRC+10011, sit apart from the rest of the sample, having the largest UD radius and monochromatic luminosity. For the other stars, there is a trend of increasing mass-loss rate with increasing UD radius and increasing monochromatic luminosity. This suggests that a maximum mass-loss rate has been reached by IRC10529 and IRC+10011 (and perhaps GY Aql, see Table 14), and further increases in UD radius, monochromatic luminosity and period (Fig. 26) do not contribute to increases in mass-loss rate. Considering the semi-regular variables only, there is no correlation between the mass-loss rate or either of the UD radius or monochromatic luminosity.
@@ 5.3.2 Binary-enhanced dust formation
Gru and R Aql have the two highest mass-loss rates for stars with periods below 300 days (Table 14). For R Aql this can be explained by its declining period (Zhao-Geisler et al., 2012); in 1915 the period was days, placing it within the trend of the other longer-period stars. The mass-loss rate calculated from CO lines is an average over several hundred years so will not have altered appreciably in that time. For Gru, the gas mass-loss rate calculated by Doan et al. (2017) takes the complex circumstellar structure into account. The high mass-loss rate, compared with other stars with pulsation periods around 200 days, can be explained by the extra dust produced in the wake of the companion, which orbits with a semi-major axis of au (approx. 3–4 stellar radii, Montargès et al. 2025; Esseldeurs et al. 2025). Once more dust is formed, radiation pressure from the AGB star will push it outwards with the rest of the dust. If it is within the wind acceleration region, this additional dust can enhance the mass loss as part of the dust-driven wind, since there will be more dust available to collide with and drag the gas outwards.
@@
The 3D single star models of Freytag & Höfner (2023) show that some gas still falls back onto the star out to radii of au, so additional dust produced within 9 au could contribute to enhanced mass-loss. This phenomenon could explain moderate mass-loss rates for the majority of surveyed SR stars with shorter periods (¡300 days). Because the shocks induced by the companion persist further out in the wind (e.g. in spiral patterns, see Appendix H and e.g. Kim et al. 2017, 2019, Malfait et al. 2021), some dust will also form well outside of the wind acceleration region (see the middle panel of Fig. 27), and will increase the overall dust-to-gas ratio compared with that of an otherwise similar single star. This could also contribute to the over-estimation of mass-loss rates calculated from dust properties and assumed dust-to-gas ratios, and could explain the discrepancy between mass-loss rates calculated from dust and from CO for the dustiest stars (e.g. see comparisons in Table 4 of Justtanont et al., 2006, where mass-loss rates from the two methods differ by around an order of magnitude).
@@
Many surveys, including both ATOMIUM and SUCCESS, are biased towards stars that have previously been detected (strongly) in CO, meaning they are likely to have higher mass-loss rates. If single stars with short pulsation periods have low levels of dust production and low mass-loss rates (e.g. McDonald et al., 2018) they would be less likely to be included in biased surveys. This explains the apparently high prevalence of binary AGB stars judging by their CO morphologies in high resolution ALMA observations (see also Sect. 5.4). An unbiased survey such as the Nearby Evolved Star Survey (NESS, Scicluna et al., 2022) is hence likely to include stars with much lower mass-loss rates than stars that have been more frequently studied for their brighter emission lines. Another example is the small fraction of dusty field stars with luminosities below the RGB tip (McDonald et al., 2012, 2017). These have been assumed to be the lowest-mass AGB stars, since they have similar luminosities to the lowest-luminosity dusty AGB stars in globular clusters (Boyer et al., 2009; McDonald et al., 2009, 2011). However, these could instead be stars that have dust production triggered by companions. The corollary to this would be that the AGB stars with the lowest mass-loss rates (perhaps or lower) are least likely to have close companions.
@@
At , around half of stars are expected to have companions with masses at least 10% that of the primary (Moe & Di Stefano, 2017). The proportion of stars with such companions increases to over 100% for stars of , meaning that some such stars are expected to be in triple systems. Close companions are more likely for higher-mass stars (Moe & Di Stefano, 2017). However, the closer the companion, the more intense the shockwaves generated as it travels through the wind and hence the more dust formation triggered, in addition to the pulsation-driven dust formation. This additional dust can enshroud the companion, making it more difficult to detect with optical telescopes. W Aql B suffers from mag of extinction au from the AGB star (Danilovich et al., 2015a) but can still be detected with Hubble and SPHERE (Ramstedt et al., 2011; Montargès et al., 2023). The close companion to Gru has only been detected by the dust associated with it in the ATOMIUM continuum (Fig. LABEL:fig:pi1gru and Homan et al. 2020) and in scattered polarised light (Montargès et al., 2023, 2025). The optical light (whether the companion is a white dwarf or a main sequence star) has been completely attenuated by the dust.
Gru is the second closest star in the ATOMIUM sample and the companion is only just resolved from the AGB continuum in the extended data. If the system were a little further away, or oriented less favourably, or the projected separation between the two stars a little smaller, then we would not be able to distinguish the companion from the AGB continuum flux. Therefore, we cannot rule out the presence of companions for the majority of the ATOMIUM sample just because we have not (yet) directly observed them. Unseen companions could partly explain the extended dust distributions seen for U Her, R Hya, and S Pav, as well as the less symmetric emission seen for GY Aql, R Aql, IRC10529, SV Aqr, T Mic and IRC+10011.
@@ 5.4 Other indications of binarity
Indirect evidence of AGB binarity has previously been found in resolved CO emission, including from the ATOMIUM data (Decin et al., 2020) and earlier studies of other stars such as R Scl (Maercker et al., 2012), CW Leo (Cernicharo et al., 2015) and AFGL 3068 (Kim et al., 2017). Such structures can be reproduced — or partially reproduced — through hydrodynamic modelling (e.g. Mastrodemos & Morris, 1998, 1999; Kim et al., 2019; El Mellah et al., 2020; Malfait et al., 2021). Molecular signatures can also be used to indirectly infer the presence of UV-emitting binaries (Van de Sande & Millar, 2022; Siebert et al., 2022; Danilovich et al., 2024). Studies surveying AGB stars in the UV or X-ray have detected emission for some stars (e.g. Ramstedt et al., 2012; Sahai et al., 2015; Ortiz & Guerrero, 2021; Schmitt et al., 2024) which ought not to come from the AGB star itself. This emission is thought to arise from companion stars, most likely from accretion discs that have formed around the companions (Sahai et al., 2015). None of the stars in the ATOMIUM sample are included in the aforementioned studies as confirmed detections151515Schmitt et al. (2024) include three ATOMIUM stars ( Gru, R Hya and S Pav) in their list of surveyed stars, but note optical contamination, which cannot be disentangled from potential X-ray detections, making their results for these stars ambiguous..
@@
Out of the 14 AGB stars in the ATOMIUM sample, we have suggested that 11 of them might have binary companions, 9 of which have not been directly detected. Given that the ATOMIUM sample was selected to cover a range of mass-loss rates and pulsation behaviours (Gottlieb et al., 2022), there was no particular expectation for the stars to have binary companions. Population statistics indicate that at least half of stars in the AGB mass range should have companions (Moe & Di Stefano, 2017), so earlier assumptions of the majority of AGB stars being single stars (i.e. before the advent of resolved imaging of the circumstellar environments) are statistically unlikely.
@@
Recent data and developments from Gaia have allowed us to better constrain distances and infer the presence of companions (including some which can be resolved Kervella et al., 2022). Based on the results of Kervella et al. (2022), Montargès et al. (2023) used a proper motion analysis combining Gaia Early Data Release 3 and Hipparcos data to estimate the probability of detecting close companions around the AGB stars. Note that the proper motion anomaly is not sensitive to wide binaries. They determine that GY Aql, R Aql, Gru, and R Hya, are likely to have a detectable close companion based on the large S/N of the proper motion anomaly, in agreement with our predictions.
@@ 6 Summary and conclusions
We present the continuum data observed for a sample of 14 AGB and 3 RSG stars as part of the ATOMIUM ALMA Large Programme. For each star, we show the continuum observations from each individual array configuration as well as the combined image. Different features are apparent on different scales, owing to resolution, maximum recoverable scales and noise levels. We identify stars with extended continuum emission on scales of 10s to 100s of au, elongated structures and a few unique structures. Notably, we see a tail of dust in the wake of the close companion to Gru and some dust forming in the wake of the distant companion to W Aql. In addition to the main ALMA array data, we present new observations taken at lower frequencies and lower resolution with the ACA, for a subsample of 7 AGB stars.
For each star, we fit uniform discs to the compact emission at the continuum peaks, which is dominated by the stellar flux but, for some stars, is likely to include contributions from dust located close to the stellar surface. We present residual images after subtracting the uniform discs from the combined continuum maps, many of which better highlight the emission offset from the continuum peaks.
We compare the uniform disc sizes and flux densities with stellar properties such as pulsation periods and mass-loss rates. Among the AGB stars, we find that all the semiregular variables in our sample have lower monochromatic luminosities and smaller UD radii than the Mira variables. As expected, the RSGs have even larger radii and higher monochromatic luminosities than the Mira variables. There is an overall correlation between the UD radii and monochromatic luminosities, as is expected from the general relationship between luminosity and radius.
@@
Comparing the ATOMIUM sample to the larger SUCCESS survey, we find that among semiregular stars with low periods, Gru has an especially large mass-loss rate, and conclude this is likely because of the additional dust forming in the wake of the shocks generated by the passage of its close companion. Hence, we propose two channels for dust formation around AGB stars: in the wake of shocks caused by pulsations and in the wake of shocks caused by the supersonic passage of a companion through the circumstellar envelope. Consequently, we propose that dust emission at mm wavelengths, offset from the continuum peak and showing coherent structure, e.g. reminiscent of a wake or shock, could be an indirect indicator of potential binarity. Deeper observations, which can detect low surface-brightness features, for more stars with known binary companions are needed to confirm this hypothesis.
Data availability
The observational data used here are openly available through the ALMA Science Archive: https://almascience.nrao.edu/aq/. The images, other data products and scripts for the standard data processing are available from the ATOMIUM Landing Page in the ALMA Science Archive: https://almascience.eso.org/alma-data/lp/atomium. The fully aligned images and other custom ALMA data products that were produced for this study are available at: doi.org/10.5281/zenodo.17015654.